The stock of “Orion S.A.” is trading €100 per share. Currently, the share capital of the
company consists of 10,000 shares and it not have any debt. Below, you can see the
balance sheet of the company in market values:
ORION Balance Sheet
Assets €1,000,000 Equity €1,000,000

“ORION S.A.” is thinking of adopting a new project that will have in present value
terms €210,000 net cash flows. The initial investment outlay for the project is only
€110,000. “Orion S.A.” is considering raising the necessary capital for this investment
by issuing new equity. All potential new shareholders will be fully aware of the
project’s value and cost, and are willing to pay fair value for the new common shares.
a. How many shares of common stock must be issued, and at what price, to raise the
required capital?
b. What is the effect, in any, of this new project on the value of the stock of the
existing shareholders?
Part B
Discuss upon the usage of the net present value rule in order to analyze mutually
exclusive projects in the cases of:
a) Postponing the investment expenditure,
b) How to choose between projects with unequal lives,
c) When to replace equipment? (<250 words)

All the rest was answered and Part B c was excluded
Answer #1

PART A

a. Shares of common stock to be issued at FAIR MARKET VALUE/PRICE = 100

No. of Shares to be issued = Initial Cash outlay / Fair Market Price = 110000 / 100 = 1100

b. Effect of a new project on the value of the Stock of the Existing Shareholders:

No of Shares of Common Stock = 1000000 / 100 + 1100 = 11100

Net Present Value of new project = 210000

Increase in Value of Shares of Common Stock by 210000.

Hence, Value of per Share of Common Stock is increased by 18.9189.

Value of Stock of Existing Shareholder = 100+18.9189 = 118.9189.

PART B

Net Present Value = Net Present Value of Cash inflow (-) Net Present Value of cash outflow

NPV RULES:

NPV >= 0, ACCEPT

NPV < 0, REJECT

a) Postponing the Investment Expenditure

If we have a proposal for investment. By using the NPV method we have to calculate the present value of the net return from the investment and if proposal not satisfied the rule of NPV method i.e. a net present value from the investment is less than zero, then we reject the proposal.

b) How to choose Between projects With unequal Lives:

when need to choose between projects of unequal lives we use NPV method with the Equivalent-Annual-Annuity Approach. We choose the project which has more Equivalent-Annualized value.

c) when to Replace Equipment:

When we compare two equipment, existing and proposed. We replace the existing equipment with the proposed equipment when the net present value of return from the proposed equipment is more.